The concentration of barium hydroxide solution is 68.4 g/dm
1
Expert's answer
2021-08-25T06:37:30-0400

M (Ba(OH)2 = 171.34 g/mol

CM = n/V

n = m/M

n (Ba(OH)2) = 68.4/171.34 = 0.4 mol

The concentration of barium hydroxide solution is 68.4 g/dm or 0.4 mol/L.